Newton is out on his wind surfer that is made to go on ice. The coefficient of static friction is 0.2 and the coefficient of kinetic friction is 0.08. The total mass of him and the windsurfer is 120kg.
A) determine the force of static friction
B) determine the force of kinetic friction
C) The wind provides a force of 250N what is his acceleration when he first starts?
D) What is his acceleration after he gets moving?
1
Expert's answer
2019-10-28T10:58:08-0400

(A) Indefinite


"F_s=\mu_s N=\mu_s mg=0.2\times 120\times 9.8=235.2\:\rm N"

(B) The force of kinetic friction  


"F_k=\mu_kN=\mu_k mg=0.08\times 120\times 9.8=94.08\:\rm N"

(C) The acceleration


"a=\frac{F}{m}=\frac{250-235.2}{120}=0.123\:\rm m/s^2"

(D)


"a=\frac{F}{m}=\frac{250-94.08}{120}=1.3\:\rm m/s^2"